OpenStudy (anonymous):

x and y intercept of ((x+1)^3 / (x-1)^2) Please!...i know that x-intercept is x=0 and y-intercept is f(0)....y-intercept result is 1 (i think)...now I have to look for the x-intercept...this is where I need help :/

OpenStudy (anonymous):

\[f(x) = ((x+1)^3 / (x-1)^2) = 0\] f(x) is undefined when : \[(x−1)^2 = 0 <=> x=1\] thus: \[(x+1)^3 = 0, x \neq1\] \[x=-1\]

OpenStudy (anonymous):

another way of thinking of it...instead of thinking of x and y intercepts...you are finding out where the equation crosses with the lines y=0(is the line that is the x axis) and x=0(which makes a line on the y axis) by plugging in those values into the equation..you are simpliy figuring out where it touches on those lines

OpenStudy (anonymous):

so the x-intercept is x=-1?

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

actually when talking about an intercept, we're talking about a point. So the intercept is (-1, 0).

OpenStudy (anonymous):

so, I only have to look the value where (in this case) the denominator = 0...so that value is the x-intercept?

OpenStudy (anonymous):

the value that make the denominator = 0 is the x-intercept?

OpenStudy (anonymous):

no. you just need to check that the denominator isn't 0 so you can just look at the numerator.

OpenStudy (anonymous):

as long as the numerator is 0, the whole fraction is 0, unless the denominator is 0 in which case the fraction doesn't exist.

OpenStudy (anonymous):

0/x = 0, x != 0

OpenStudy (anonymous):

ok...and in this exercise?? e^-x sen(x)

OpenStudy (anonymous):

(e^-x)(sen x)...find the x and y intercept

OpenStudy (anonymous):

I have that the y-intercept is 0...but I'm confused about the x-intercept

OpenStudy (anonymous):

you get two equations. e^-x = 0 or sin(x) = 0

OpenStudy (anonymous):

then combine the answers.

OpenStudy (anonymous):

what do you mean by combine the answers?

OpenStudy (anonymous):

I know that sin(x)= 0 on every pi

OpenStudy (anonymous):

but, when e^-x=0??

OpenStudy (anonymous):

well from sin(x) = 0 you get PI*n, n is Z and from e^-x you don't get a zero

OpenStudy (anonymous):

so..it doesn't have x-intercepts?

OpenStudy (anonymous):

the answer is \[\pi*n, n \in Z\]

OpenStudy (anonymous):

no, e^-x doesn't have x-intercepts

OpenStudy (anonymous):

wait...you told me that the answer is pi n and then that it doesn't have x-intercepts :/

OpenStudy (anonymous):

the problem is (e^-x)(sen x)

OpenStudy (anonymous):

to the whole problem the answer is PI n

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

this is called the zero-product property

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

could yo verify this intercepts please?

OpenStudy (anonymous):

a. (x^2-16)/(x^2-4x) [x=>-4 and 4][y=>0] b. (x)/(x^2 -9) [x=>0][y=>0] c. (x)/(squared root x^2 +6) [x=>0][y=>0] d. (sin x)/(1+cos x) [x=>pi n][y=>0] e. (e^-x)(sen x) [x=>pi n][y=>0] f. (x+1)^3 / (x-1)^2 [x=>-1][y=>1]

OpenStudy (anonymous):

you can verify these yourself. just plug the answer to the expression.

OpenStudy (anonymous):

what result do I have to get so i Know that my result is right?

OpenStudy (anonymous):

if you plug the x-intercept, you need to get 0 as the result. if you plug the y-intercept (x=0), you need to get the answer as the result.

OpenStudy (anonymous):

thank you A LOT!